Use the quadratic formula to solve each of the following quadratic equations.
3x^2-2x+1=0

Answer by checkley75(3666)

3X^2-2X+1=0
X=(2+-SQRT[-2^2-4*3*1])/2*3
X=(2+-SQRT4-12])/6
X=(2+-SQRT-8)/6
X=(2+-2.828i)/6
X=(2+2.828i)/6
X=2/6+2.828i/6
X=1/3+.4714i ANSWER.
X=(2-2.828i)/6
X=1/3-.4714i ANSWER.